# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

### v2

@@ IV Numerical Results
In this section, we present numerical results to validate our proposals. The experimental setup is illustrated in Fig. 2, featuring multiple TBS and RBS units jointly attempting to detect a target amidst clutters while simultaneously respectively serving two DL and two UL users. In the experiment, TBSs, RBSs, DL users, UL users, one target and clutters are randomly located within a circular area with a radius of $70$ meters (m). We employ a geometric channel model [4] for the link concerning each user, wherein the number of transmit and receive channel paths is consistently identical, i.e., $L^{t,0}_{m,k}=L^{r,0}_{m,k}=L^{t,1}_{p,l}=L^{r,1}_{p,l}=L^{t,2}_{k,l}=L^{r,2}_{k,l}\triangleq L,\forall m\in\mathcal{M}_{t},\forall p\in\mathcal{M}_{r},\forall k\in\mathcal{K}_{d},\forall l\in\mathcal{K}_{u}$. Thus, the path-response matrix for each user is diagonal, i.e., the path-response matrix for between the $m$-th TBS and the $k$-th DL user is represented as $\mathbf{\Sigma}^{0}_{m,k}=\textrm{diag}\{\sigma^{0}_{m,k,1},\dots,\sigma^{0}_{m,k,L}\}$, where each $\mathbf{\sigma}^{0}_{m,k,l}$ satisfying $\sigma^{0}_{m,k,l}\sim\mathcal{CN}(0,{(c^{0}_{m,k})^{2}}/{L}),l=1,\dots,L$ [4], and $(c^{0}_{m,k})^{2}=C_{0}d_{m,k}^{-\alpha_{loss}}$, where $C_{0}$ corresponds to the path loss at the reference distance of $1$ m, $d_{m,k}$ is the propagation distance between the $m$-th TBS and the $k$-th DL user, $\alpha_{loss}=2.8$ is the path-loss exponent. The elevation and azimuth of both AoAs and AoDs are assumed to be independent and identically distributed variables following the uniform distribution over $[-\frac{\pi}{2},\frac{\pi}{2}]$. The TBS-target/clutters and RBS-target/clutters are modeled as line-of-sight (LoS) channels. The moving regions for MAs at both the BS and users are set as square areas of size $[-\frac{A}{2},\frac{A}{2}]\times[-\frac{A}{2},\frac{A}{2}]$. In addition, $M_{t}=2$ TBS equip $N_{t}=4$ antennas and $M_{r}=2$ RBS equip $N_{r}=4$ antennas. The transmit power of each TBS is set as $30$dBm. The noise power and the predefined target detection level of RBS are set as $\sigma^{2}_{d,k}=\sigma^{2}_{r}=-80$dBm and $\Gamma_{r}=3$dB, respectively.
@@
Fig. 3 examines the convergence behaviour of the overall solution in Alg. 1. The obtained sun-rate iterates under varying configurations of $N_{t}$ and $N_{r}$ are illustrated in the plot. As shown in Fig. 3, the proposed algorithm can yield monotonic improvement in sun-rate and generally achieves significant beamforming gain within the first 10 iterations.
@@
Alg. 1 is labeled by “Joint MA”. For comparison in sum-rate enhancement, we consider the following three foundational scenarios: 1) BS MA: All BSs are equipped with the MAs, while all users respectively employ an MA fixing at the original point. 2) User MA: All users respectively employ an MA, while the MAs of all BSs are fixed. 3) FPA: Both all BSs and users are equipped with FPA-based array.
Fig. 6 compares the sum-rate for different schemes versus the transmit power of all TBS. The sum-rate of all schemes increases significantly with the transmit power and the three MA-based schemes significantly improve the sum-rate compared to the FPA-based case. In particular, the “Joint MA” case shows a notable advantage over all other cases in terms of sum-rate, demonstrating performance improvement of approximately 20.0% over the FPA scheme. This is anticipated, as the proposed “Joint MA” scheme exploits the most spatial DoFs, and the other three schemes face performance loss due to less flexibility in channel reconfiguration.
Fig. 6 illustrates the impact of the number of antenna at the TBSs. The left and right subplots corresponding to different numbers of TBSs, respectively. In our test, the number of antenna $N_{t}$ varies from 2 to 9. As reflected by Fig. 6, increasing $N_{t}$ can improve beamforming gain for all schemes. Moreover, the “Joint MA” scheme boosts the system’s sum-rate more significantly compared to all other schemes. Fig. 6 shows the impact of the number of antenna at the RBSs. As can be seen in both left and right subplots, which corresponding to varying quantities of RBSs, sum-rate increases when $N_{r}$ increases. Compared to other foundational scenarios, the deployment of MA in both BS and user can significantly boot the sum-rate.
Figure 7: The impact of the number of paths for DL user.
Figure 8: The impact of the number of paths for UL user.
Figure 9: The impact of the normalized region size.
@@
Fig. 9 describes the sum-rate versus the normalized region size $A/\lambda$ for MA at the BSs. As can be observed, both the “Joint MA” and “BS MA” schemes experience increases as the normalized region size grows, attributed to the larger moving area of BSs’ MA providing more spatial DoFs. Besides, we observe that the increase of the normalized region size will enlarge the performance gaps between “Joint MA” and “BS MA” schemes. Since both BSs and users are equipped with the MA, the “Joint MA” case can provide more effective channel reconfiguration.
@@ V Conclusions
This paper considers the joint active beamforming and position coefficients design problem in an MA aided networked FD ISAC system that accomplishes radar sensing as well as DL and UL communication capabilities concurrently. We propose an iterative algorithm to jointly optimize BS beamforming, UL users’ power allocation, receiving processors and MA configuration for enabling radar sensing and communication functionalities. Numerical results showcase the effectiveness of our proposed solution and highlight the advantages of deploying MA in the networked FD ISAC system.
-A Derivations of $\nabla h^{R1}_{k,m,n,i,j}$ and $\psi_{1,k,m,n,i,j}$
Firstly, the elements of $\nabla h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})=\big{[}\frac{\partial\nabla h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})}{\partial x^{d}_{k,0}},\frac{\partial\nabla h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})}{\partial y^{d}_{k,0}}\big{]}^{T}$ are given as
$\frac{\partial h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}}\!\!\!=\!\!-\!\frac{2\pi}{\lambda}\!|\!\mathbf{Q}^{1}_{k,m,n}\![i,j]\!|\!\big{(}\!\!\!-\!\cos(\!\theta^{r,0}_{m,k,i}\!)\!\sin(\!\phi^{r,0}_{m,k,i}\!)$ (94)
$\!\!+\!\!\cos(\!\theta^{r,0}_{n,k,j}\!)\!\sin(\!\phi^{r,0}_{n,k,j}\!)\!\big{)}\!\!\sin\!\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{\!T}\!\!(\!\mathbf{a}^{r,0}_{n,k,j}\!\!\!-\!\mathbf{a}^{r,0}_{m,k,i}\!)\!\!+\!\!\!\angle\!\mathbf{Q}^{1}_{k,m,n}\![i,\!j]\!\big{)},$
$\frac{\partial h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})}{\partial y^{d}_{k,0}}=-\!\frac{2\pi}{\lambda}|\mathbf{Q}^{1}_{k,m,n}[i,j]|(-\sin(\theta^{r,0}_{m,k,i})$ (95)
$+\sin(\theta^{r,0}_{n,k,j}))\sin\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{T}\!(\!-\!\mathbf{a}^{r,0}_{m,k,i}\!+\!\mathbf{a}^{r,0}_{n,k,j})\!+\!\angle\!\mathbf{Q}^{1}_{k,m,n}[i,j]\big{)}.$
To obtain $\psi_{1,k,m,n,i,j}$, the Hessian matrix of $\nabla^{2}h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})$ over $\mathbf{r}_{0,k,0}$ can be formulated as
$\nabla^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)\!\!\triangleq\!\!\!\!\begin{bmatrix}\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial x^{d}_{k,0}}&\!\!\!\!\!\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial y^{d}_{k,0}}\\ \frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial x^{d}_{k,0}}&\!\!\!\!\!\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial y^{d}_{k,0}}\end{bmatrix},$ (96)
where the matrix units are respectively given as follows
$\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}\!(\!\mathbf{r}_{0,k,0}\!\!)}{\partial x^{d}_{k,0}\partial x^{d}_{k,0}}\!\!\!=\!\!-\!\frac{4\pi^{2}}{\lambda^{2}}|\!\mathbf{Q}^{1}_{k,m,n}[i,j]\!|\!\big{(}\!\!\!-\!\!\cos\!(\!\theta^{r,0}_{m,k,i}\!)\!\sin(\!\phi^{r,0}_{m,k,i}\!)$ (97)
$\!\!+\!\!\cos(\!\theta^{r,0}_{n,k,j}\!)\!\sin(\!\phi^{r,0}_{n,k,j}\!)\!\big{)}^{\!2}\!\!\!\cos\!\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{\!T}\!\!(\!\mathbf{a}^{r,0}_{n,k,j}\!\!\!-\!\mathbf{a}^{r,0}_{m,k,i}\!)\!\!+\!\!\!\angle\!\mathbf{Q}^{1}_{k,m,n}\![i,\!j]\!\big{)},$
$\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial y^{d}_{k,0}}\!\!=\!\!\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial x^{d}_{k,0}}\!\!\!=\!\!-\!\frac{4\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{1}_{k,m,n}[i,j]|$ (98)
$\big{(}\!\!\!-\!\cos(\!\theta^{r,0}_{m,k,i}\!)\!\sin(\!\phi^{r,0}_{m,k,i}\!)\!\!+\!\!\cos(\!\theta^{r,0}_{n,k,j}\!)\sin(\!\phi^{r,0}_{n,k,j}\!)\!(-\sin(\theta^{r,0}_{m,k,i})$
$+\sin(\theta^{r,0}_{n,k,j}))\cos\!\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{T}\!(\!-\!\mathbf{a}^{r,0}_{m,k,i}\!\!+\!\mathbf{a}^{r,0}_{n,k,j}\!)\!\!+\!\!\angle\!\mathbf{Q}^{1}_{k,m,n}[i,j]\!\big{)}\big{)},$
$\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial y^{d}_{k,0}}=-\!\frac{4\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{1}_{k,m,n}[i,j]|(-\sin(\theta^{r,0}_{m,k,i})$ (99)
$+\sin(\theta^{r,0}_{n,k,j}))^{2}\cos\!\!\big{(}\!\frac{2\pi}{\lambda}\!(\!\mathbf{r}_{0,k,0}\!)^{T}\!(\!-\!\mathbf{a}^{r,0}_{m,k,i}\!\!+\!\mathbf{a}^{r,0}_{n,k,j}\!)\!\!+\!\!\angle\!\mathbf{Q}^{1}_{k,m,n}[i,j]\!\big{)}.$
And then, we have
$\|\nabla^{2}h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})\|_{2}\leq\|\nabla^{2}h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})\|_{F}$ (100)
$=\bigg{[}\bigg{(}\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial x^{d}_{k,0}}\bigg{)}^{2}+\bigg{(}\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial x^{d}_{k,0}\partial y^{d}_{k,0}}\bigg{)}^{2}$
$+\bigg{(}\frac{\partial^{2}\!h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial x^{d}_{k,0}}\bigg{)}^{2}+\bigg{(}\frac{\partial^{2}h^{R1}_{k,m,n,i,j}(\!\mathbf{r}_{0,k,0}\!)}{\partial y^{d}_{k,0}\partial y^{d}_{k,0}}\bigg{)}^{2}\bigg{]}^{\frac{1}{2}}$
$\leq\bigg{[}4\bigg{(}\frac{16\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{4}_{k,m,n}[i,j]\bigg{)}^{2}\bigg{]}^{\frac{1}{2}}.$
Therefore, to guarantee that operation (a) in (80) yields an upper-bound for $h^{R1}_{k,m,n,i,j}$, we can define $\psi_{1,k,m,n,i,j}=\frac{32\pi^{2}}{\lambda^{2}}|\mathbf{Q}^{1}_{k,m,n}[i,j]$ with $\psi_{1,k,m,n,i,j}\mathbf{I}_{2}\succeq\nabla^{2}h^{R1}_{k,m,n,i,j}(\mathbf{r}_{0,k,0})$.

### v3

@@ IV Numerical Results
In this section, we present numerical results to validate our proposals. The experimental setup is illustrated in Fig. 2, featuring multiple TBS and RBS units jointly attempting to detect a target amidst clutters while simultaneously respectively serving three DL and three UL users. In the experiment, TBSs, RBSs, DL users, UL users, one target, and clutters are randomly located within a circular area with a radius of $70$ meters (m). We employ a geometric channel model [5] for the link concerning each user, wherein the number of transmit and receive channel paths is consistently identical, i.e., $L^{t,0}_{m,k}=L^{r,0}_{m,k}=L^{t,1}_{p,l}=L^{r,1}_{p,l}=L^{t,2}_{k,l}=L^{r,2}_{k,l}\triangleq L=6,\forall m\in\mathcal{M}_{t},\forall p\in\mathcal{M}_{r},\forall k\in\mathcal{K}_{d},\forall l\in\mathcal{K}_{u}$.Thus, the path-response matrix for each user is diagonal, i.e., the path-response matrix for between the $m$-th TBS and the $k$-th DL user is represented as $\mathbf{\Sigma}^{0}_{m,k}=\textrm{diag}\{\sigma^{0}_{m,k,1},\dots,\sigma^{0}_{m,k,L}\}$, where each $\mathbf{\sigma}^{0}_{m,k,l}$ satisfying $\sigma^{0}_{m,k,l}\sim\mathcal{CN}(0,{(c^{0}_{m,k})^{2}}/{L}),l=1,\dots,L$ [5], and $(c^{0}_{m,k})^{2}=C_{0}d_{m,k}^{-\alpha_{loss}}$, where $C_{0}$ corresponds to the path loss at the reference distance of $1$ m, $d_{m,k}$ is the propagation distance between the $m$-th TBS and the $k$-th DL user, $\alpha_{loss}=2.8$ is the path-loss exponent. The elevation and azimuth of both AoAs and AoDs are assumed to be independent and identically distributed variables following the uniform distribution over $\frac{\pi}{2},\frac{\pi}{2}]$. The TBS-target/clutters and RBS-target/clutters are modeled as line-of-sight (LoS) channels. The moving regions for MAs at both the BS and users are set as square areas of size $\frac{A}{2},\frac{A}{2}]\times[-\frac{A}{2},\frac{A}{2}]$. In addition, $M_{t}=2$ TBS equip $N_{t}=4$ MAs and $M_{r}=2$ RBS equip $N_{r}=4$ MAs. The transmit power of each TBS is set as 3030dBm. The noise power and the predefined target detection level of RBS are set as $\sigma^{2}_{d,k}=\sigma^{2}_{r}=-80$dBm and $\Gamma_{r}=3$dB, respectively.
@@
Fig. 3 examines the convergence behavior of the overall solution in Alg. 1. The obtained sum-rate iterates under varying configurations of $N_{t}$ and $N_{r}$ are illustrated in the plot. As shown in Fig. 3, the proposed algorithm can yield monotonic improvement in sum-rate and generally achieves significant beamforming gain within the first 10 iterations.
@@
Alg. 1 is labeled by “Joint MA”. For comparison in sum-rate enhancement, we consider the following four foundational scenarios: 1) BS MA: All BSs are equipped with the MAs, while all users respectively employ an MA with fixed; 2) User MA: All users respectively employ an MA, while the MAs of all BSs are fixed; 3) Rand. MA: The MAs of all BSs and users have been fixed; 4) FPA: Both BSs and users are equipped with the FPA-based array.
Fig. 4 compares the sum-rate for different schemes versus the transmit power of all TBS. The sum-rate of all schemes increases significantly with the transmit power and the three MA-based schemes, i.e., “Joint MA”, “BS MA”, and “User MA” cases, significantly improve the sum-rate compared to both “FPA” and “Rand. MA” cases. For the “Joint MA” case, when the transmit power increases from 10 dBm to 35 dBm, the sum-rate increases by about 67%. In particular, the “Joint MA” case shows a notable advantage over all other cases regarding sum-rate. Specifically, under the 20 dBm case, the “Joint MA” case outperforms “FPA” and “Rand. MA” cases by approximately 30% and 35%, respectively. This is anticipated, as the proposed “Joint MA” scheme exploits the most spatial DoFs, and the other four schemes face performance loss due to less flexibility in channel reconfiguration.
@@
Fig. 5 illustrates the impact of the number of antenna at the TBSs. The upper and lower subplots correspond to different numbers of TBSs, respectively. In our test, the number of antenna $N_{t}$ varies from 2 to 9. As reflected by Fig. 5, increasing $N_{t}$ can improve beamforming gain for all schemes. Moreover, the “Joint MA” scheme boosts the system’s sum-rate more significantly compared to all other schemes. For the “Joint MA” case, when the transmit antenna number increases from 2 to 9, the sum-rate improves by approximately 25% for the $M_{t}=2$” scenario and by about 33% for the $M_{t}=4$” scenario.
Fig. 6 shows the impact of the number of antenna at the RBSs. As can be seen in both upper and lower subplots, which correspond to varying quantities of RBSs, sum-rate increases when $N_{r}$ increases. Compared to other benchmarks, the deployment of MA in both BS and user can significantly boost the sum-rate.
Figure 7: The impact of the number of paths between TBS and DL user.
Figure 8: The impact of the number of paths between RBS and UL user.
@@
Figure 9: The impact of the sensing SINR performance.
The impact of varying sensing SINR requirements is illustrated in Fig. 9. A clear trade-off between communication performance and target sensing can be observed. In this experiment, the sensing SINR ranges from 0 dB to $5$ dB. As shown in the figure, increasing the sensing SINR results in a decrease in the sum-rate for all cases.
@@ V Conclusions
In this paper, we considered a novel MA-aided networked ISAC system that accomplishes radar sensing as well as DL and UL communication capabilities concurrently. Aiming to maximize the sum-rate among all DL and UL users, we investigated a joint active beamforming and antenna position coefficients design problem. Despite the non-convex nature of the considered problem with highly coupled optimization variables, we propose an iterative algorithm to jointly optimize BS beamforming, UL users’ power allocation, receiving processors and MAs’ position configuration for enabling radar sensing and communication functionalities. Numerical results showcase the effectiveness of our proposed solution, highlight the advantages of deploying MA in the networked ISAC system, and offer useful engineering insights. Firstly, while ensuring the sensing performance, MA can enhance communication performance compared to the case with the FPA-based array. Additionally, in comparison to benchmark schemes, our proposed algorithm significantly reduces the required transmit power or the number of antennas needed to achieve the desired level of communication performance.